Answer the following questions related to the given electrochemical cell.​Answer the following questions related to the given electrochemical cell 2NO(g) H20(l) 2e N20(g) 20H (aq) 0.760 V NO aq) H2O(l) 2e NO2 (aq) 20H (aq) Eo 0.010 V 1. Answer the following questions under standard conditions (a) The half cell containing N20/NO is the (b) The half cell containing NO3 /NO2 s the (c) What is Eocell (in V)? Report your answer to three decimal places in standard notation (i.e., 0.123 V). Submit Answer Tries 0/3
Answer #1

(1)

(a) Cathode half cell - reduction half cell

(b) Anode half cell - oxidation half cell

(c) E0cell = E0cathode - E0anode = 0.760 - 0.010 = 0.750 V

(2)

(b) E = E0 + (RT/nF) Log(PNO/{PN2O*[OH-]})

(3)

(d) E = E0 + (RT/nF) LOg[NO3-]a/[NO2]c[OH-]d